OpenStudy (anonymous):

how do you solve 3x^2-5=0 using the quadratic formula?

OpenStudy (campbell_st):

the quadratic formula is used for a quadratic in the form \[ax^2 + bx + c = 0\] and the GQF is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a = 3, b = 0 and c = -5 substitute and evaluate